Homework: HW-6: Chapter 7 Score: 0 of 1 pt 6 of 10 (6 complete) X Text Question 2.13 Assume the short run variable cost functThe total cost of production (C) is The marginal cost of production (MC) is The average fixed cost of production (AFC) is TheHW Score: 48%, 4.8 of 10 pts 6 of 10 (6 complete) EQuestion Help VC =05q0.8 (VC), the marginal cost of production (MC), the aAssume the short run variable cost function for Japanese beer is VCequals0.5q Superscript 0.8. If the fixed cost​ (F) is ​$600 and the firm produces 400 ​units, determine the total cost of production​ (C), the variable cost of production​ (VC), the marginal cost of production​ (MC), the average fixed cost of production​ (AFC), and the average variable cost of production​ (AVC). What happens to these costs if the firm increases its output to 500​? Assuming the firm produces 400 ​units, the variable cost of production​ (VC) is VCequals nothing. ​(Enter your response rounded to two decimal​ places.)

Homework: HW-6: Chapter 7 Score: 0 of 1 pt 6 of 10 (6 complete) X Text Question 2.13 Assume the short run variable cost function for Japanese beer is VC=0.5q08 If the fixed cost (F) is $600 and the firm produces 400 units, determine the total cost of production (C), the variable cost of production (VC), the marginal cost of production (MC), the average fixed cost of prod output to 500? Assuming the firm produces 400 units, the variable cost of production (VC) is (Enter your response rounded i two decimal places) VC Enter your answer in the answer box and then click Check Answer Clear All paris ining
The total cost of production (C) is The marginal cost of production (MC) is The average fixed cost of production (AFC) is The average variable cost of production (AVC) is Now suppose the firm increases output to 750 units. The variable cost of production (VC) is The total cost of production (C) is The marginal cost of production (MC) is The average fixed cost of production (AFC) is The average variable cost of production (AVC) is Question is complete. Tap on the red indicators to see incorrect answers. All parts showing
HW Score: 48%, 4.8 of 10 pts 6 of 10 (6 complete) EQuestion Help VC =05q0.8 (VC), the marginal cost of production (MC), the average fixed cost of production (AFC), and the average variable cost of production (AVC). What happens to these costs if the firm increases its two decimal places.) (Enter your response rounded VC ? Check Anser Clear All 23:56 英M 42 2019/6/16
Answer #1


Variable Cost ( VC ) (q)=0.5*q^0.8

FC=$600
q=400 units

Total Cost=Fixed Cost+Variable Cost=600+0.5(400^0.8)=$660.341

Variable Cost=TC-FC=660.341-600=60.341

Marginal Cost=dTC/dq=0.8*0.5*q^(0.8-1)=0.4*q^-0.2

MC at q=400 is MC=0.4*(400)^-0.2=0.12068

AFC=FC/q=600/q=600/400=1.5

AVC=VC/q=0.5*q^0.8/q=0.5*q^-0.2=0.5(600)^-0.2=0.1391


Now q=750 units


VC=0.5*q^0.8=0.5*(750)^0.8=99.774

TC=FC+VC=600+99.77=699.77

MC=0.4*q^-0.2=0.4*(750)^-0.2=0.104

AFC=FC/q=600/q=600/750=$0.8

AVC=VC/q=0.5q^0.8/q=0.5q^-0.2=0.5(750)^-0.2=0.133


Now q=500 units

VC=0.5*q^0.8=0.5*(500)^0.8=72.13

TC=FC+VC=600+72.13=672.13

MC=0.4*q^-0.2=0.4*(500)^-0.2=0.1154

AFC=FC/q=600/q=600/500=$1.2

AVC=VC/q=0.5*q^0.8/q=0.5*q^-0.2=0.5(500)^-0.2=0.1442